OpenStudy (anonymous):

x/3 + x/4 = 1 Solve and round to the nearest tenth

OpenStudy (anonymous):

4x+3x/12 = 1 7x = 12 x = 12/7 = 1.7142857 Then, round it off to nearest tenth. 1.70

OpenStudy (anonymous):

Lol Thank you.